OpenStudy (jiteshmeghwal9):

@ParthKohli @hartnn first of all give me a revision on quadratic equations & give me some tips on them so i can write notes on them & then tell me a site where i can gt a lot of question on quadratic equation :)

Parth (parthkohli):

Paul's Notes and Khan Academy!

OpenStudy (jiteshmeghwal9):

I know khan academy but what is the site for paul's notes ?

OpenStudy (jiteshmeghwal9):

@ParthKohli is there any google search for them

Parth (parthkohli):

http//khanacademy.org http://tutorial.math.lamar.edu

Parth (parthkohli):

http://khanacademy.org *

Parth (parthkohli):

How could I forget AoPS?

Parth (parthkohli):

Wait, take a look at the videos.

OpenStudy (jiteshmeghwal9):

well nice ntes on paul maths notes but wht is aops

Parth (parthkohli):

One second...

Parth (parthkohli):

I bet this will be enough for ya :)

OpenStudy (jiteshmeghwal9):

ya & patrickJMT.com is also best :)

Parth (parthkohli):

Yup, that guy is really great!

Parth (parthkohli):

I want his Sharpies. :P

Parth (parthkohli):

You can use AoPS Alcumus for practicing: http://www.artofproblemsolving.com/Alcumus/Introduction.php

OpenStudy (jiteshmeghwal9):

sharpies=?

Parth (parthkohli):

The pens he uses.

OpenStudy (jiteshmeghwal9):

haha , i want his pages he rubs his mistakes with his hands :)

Parth (parthkohli):

Haha yes. :p

OpenStudy (jiteshmeghwal9):

^_^

Parth (parthkohli):

Just take a look at AoPS videos -- the man who did the videos also sometimes, dresses as a black sweater guy who uses tricks to solve problems.

OpenStudy (jiteshmeghwal9):

ok!

OpenStudy (hba):

Well , The first step is to convert it to the standard form of Quadratic equation. Which is, \[\huge\ ax^2+bx+c=0\] You should also keep in mind the Quad formula. \[\huge\ x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] There are some other methods to solve the Quad equations. Which are mentioned below, 1)Breaking the middle term. 2)Completing the whole square method.

Parth (parthkohli):

Do you need the derivation of Quadratic Formula? @jiteshmeghwal9 =)

Parth (parthkohli):

I'd just write it down.\[ax^2 + bx + c = 0\]Multiply \(4a\) to both sides.\[4a^2x^2 + 4abx + 4ac = 0\]Subtract \(4ac\) from both sides.\[4a^2x^2+ 4ab = -4ac\]Add \(b^2\) to both sides.\[4a^2x^2 + 4ab +b^2= b^2 - 4ac\]Now completing the square,\[(2ax + b)^2 = b^2 - 4ac\]Square root\[2ax + b = \pm\sqrt{b^2 - 4ac}\]\[2ax = -b \pm \sqrt{b^2 - 4ac}\]\[\boxed{x = \dfrac{-b\pm \sqrt{b^2 - 4ac}} {2a}} \]

OpenStudy (jiteshmeghwal9):

I know the quadratic equation's derivation. It's easy. i have a tutorial on this also :)

Parth (parthkohli):

Nice!

OpenStudy (jiteshmeghwal9):

well, paul's notes are very interesting :)

Parth (parthkohli):

They are, yes :D

OpenStudy (jiteshmeghwal9):

when i wrote the tutorial @lgbasallote asked me so tough questions on that i totally confused on that so i asked u for the notes

OpenStudy (jiteshmeghwal9):

now i'm going log out & the viper is coming online, bye :)